# Akaki Mamageishvili Offchain Labs Benny Sudakov ETH Zurich

### v1

@@ 3.2 PoS Savings
Suppose a PoS protocol centered at any service $s\in S$, having security parameter $\alpha_{s}$ and allocated $\sigma_{v}^{s}$ stakes for any $v\in N_{G}(s)$ is safe, by the same definition as above: no attacking set of validators find it profitable to attack service $s$. Let the partition of the stake vector $\sigma$ be denoted by $\sigma^{S}$. Consider the total stake for any validator $v\in V$ defined as $\sigma_{v}=\sum_{s\in N_{G}(v)}$. That is, we aggregate all stakes that the validator has staked across all PoS protocols in which it participates. We ask a similar question to the previous section:
@@
First, we note the following inequality that relates the total value to capture from services and the total staked amount of validators. It will be useful to obtain an upper bound on the PoS savings.
@@
$T(\sigma)\geq T(\pi)$.
@@ A PoS protocol centered at $s\in S$ is secure implies that
Consider PoS protocols with $m$ services, each service $s\in S$ having value $\pi_{s}=1$, security parameter $\alpha_{s}=\frac{1}{m+1}$ and two validators, labeled $s$ and ${m+1}$ with stakes equal to $\sigma_{s}^{s}=1$ and $\sigma_{m+1}^{s}=\frac{1}{m}$. Note that all PoS protocols are secure:
@@
  • small staked validator ${m+1}$ can not attack the service since its stake share is equal to $\frac{\frac{1}{m}}{1+\frac{1}{m}}=\frac{1}{m+1}$, that is the same as a security parameter of the service.
We claim that the minimum extra stake required is $m-1$. In the resulting restaking protocol, validator ${m+1}$ has stake equal to $m\frac{1}{m}=1$. It can attack all services unless one of the other validators gets an extra $m-1$ stake, in which case the claim is trivially true. This amount is enough. Consider adding stake $m-1$ to the validator ${m+1}$ so that it now controls the stake amount $m$. This validator does not attack services since it is losing at least the amount she can gain by attacking. Other validators are not able to attack their corresponding services, and it is also not profitable to attack.
@@ 4 Conclusion
We started a framework to compare restaking and implied Proof-of-Stake protocols in terms of total stake requirements. We provide asymptotically matching lower and upper bounds on restaking savings, as well as a constant lower bound on the PoS savings and nonconstant upper bounds on it. Our results indicate that none of the solution concepts dominates another. Closing the gap for PoS savings remains an interesting question. Another open question is to come up with a (generic) sufficient condition for a restaking graph to be secure, that can be checked in a polynomial time, and the condition accepts some graphs with positive restaking savings.

### v2

@@ 3.2 PoS Savings
Suppose a PoS protocol centered at any service $s\in S$, having security parameter $\alpha_{s}$ and allocated $\sigma_{v}^{s}$ stakes for any $v\in N_{G}(s)$ is safe, by the same definition as above: no attacking set of validators find it profitable to attack service $s$. Let the partition of the stake vector $\sigma$ be denoted by $\sigma^{S}$. Consider the total stake for any validator $v\in V$ defined as $\sigma_{v}=\sum_{s\in N_{G}(v)}\sigma_{v}^{s}$. That is, we aggregate all stakes that the validator has staked across all PoS protocols in which it participates. We ask a similar question to the previous section:
@@
First, we note the following inequality that relates the total value to capture from all services and the total staked amount of validators. It will be useful to obtain an upper bound on the PoS savings.
@@
The total stake amount of all validators is more or equal than the total value derived from all projects: $T(\sigma)\geq T(\pi)$.
@@ A PoS protocol centered at $s\in S$ is secure implies that
Consider PoS protocols with $m$ services, each service $s\in S$ having value $\pi_{s}=1$, security parameter $\alpha_{s}=\frac{1}{m+1}$ and two validators, labeled $s$ and m+1{m+1} with stakes equal to $\sigma_{s}^{s}=1$ and $\sigma_{m+1}^{s}=\frac{1}{m}$. Note that all PoS protocols are secure:
@@
  • small staked validator m+1{m+1} can not attack the service since its stake share is equal to $\frac{\frac{1}{m}}{1+\frac{1}{m}}=\frac{1}{m+1}$, that is the same as a security parameter of the service.
We claim that the minimum additional stake required is m1m-1. In the resulting restaking protocol, the validator m+1{m+1} has stake equal to $m\frac{1}{m}=1$. It can attack all services unless one of the other validators gets an extra m1m-1 stake, in which case the claim is trivially true. This amount is enough. Consider adding stake m1m-1 to the validator m+1{m+1} so that it now controls the stake amount $m$. This validator does not attack services since it is losing at least the amount she can gain by attacking. Other validators are not able to attack their corresponding services, and it is also not profitable to attack.
@@
Interpretation of Results
Although mostly conceptual, our results can provide some design recommendations. For an existing restaking project that has an underlying graph structure already with specified stake sizes and service valuations, we could compute restaking savings. For a given restaking graph, it must be feasible to check if it is secure using heuristic algorithms, instead of checking a polynomial-time sufficient condition which severely limits the domain and may result into zero restaking savings.
The upper bound obtained in Proposition 4 is interesting in practical cases, since the security parameters of most protocols are usually from the set of fractions $\{\frac{1}{4},\frac{1}{3},\frac{1}{2}\}$. These values come from Byzantine fault-tolerant mechanisms, used to achieve consensus on the state of the chain.
The number of unique addressed validators is high for the Ethereum chain, approaching one million, while the real number of validators can still be on the order of tens of thousands. All validators are staked for Ethereum, that is, the restaking protocol based on Ethereum has the highest degree $\max_{s}{d_{G}(s)}=n$. The number of projects that require staking services is typically low, in the order of hundreds, and it is highly unlikely that a validator would provide services for most of them, therefore, it would be expected that $\max_{v}(d_{G}(v))$ is low.
Recently, EigenLayer’s restaking project upgraded the system that only allows a unique slashing property: any staked asset can be slashed by a single service  [^4] https://blog.eigencloud.xyz/introducing-the-eigenlayer-security-model/. This makes it equivalent to validators distributing their total stake amount for separate services, and hence results into separate secure PoS protocols.
@@ 4 Conclusion
We initiate a framework to compare restaking and implied Proof-of-Stake protocols in terms of total stake requirements. We provide asymptotically matching lower and upper bounds on restaking savings, as well as a constant lower bound on the PoS savings and non-constant upper bounds on it. Our results indicate that none of the solution concepts dominates another. Closing the gap for PoS savings remains an open question. Another open question is to come up with a (generic) sufficient condition for a restaking graph to be secure, that can be checked in a polynomial time, and the condition accepts some graphs with positive restaking savings.